# Inferring Dynamic Networks from Marginals with Iterative Proportional Fitting

### v1

@@ 1 Introduction
One approach to solving this problem is to repurpose the classic iterative proportional fitting (IPF) procedure (Deming & Stephan, 1940), also widely known as Sinkhorn’s algorithm (Sinkhorn, 1974). Given a matrix $X$, target row marginals $p$, and target column marginals $q$, IPF tries to find a biproportional scaling of $X$ to match the target marginals. IPF poses an attractive solution for the network inference problem: it matches the marginal constraints of the problem, is computationally lightweight and space efficient, and is supported by decades of literature devoted to its analysis (Bregman, 1967; Fienberg, 1970; Csiszár, 1975; Pukelsheim & Simeone, 2009; Léger, 2021; Qu et al., 2023). Furthermore, prior works have shown empirical success with using IPF to infer networks (McCord et al., 2010; Chang et al., 2021a), enabling important applications such as modeling COVID-19 spread on detailed mobility networks (Li et al., 2023; Chaudhuri et al., 2022; Alimohammadi et al., 2023) and supporting policymakers (Chang et al., 2021b).
@@ 2 Related Work
Iterative proportional fitting. IPF (Deming & Stephan, 1940), also known as Sinkhorn’s algorithm, biproportional fitting, raking, and the RAS algorithm, has a long history across disciplines. IPF offers many empirical advantages: it is straightforward to implement (see Algorithm 1 in Appendix A), transparent, reproducible, space efficient, and computationally lightweight (Lomax & Norman, 2015; Lovelace et al., 2015). The algorithmic properties of IPF have also been extensively studied (Sinkhorn, 1964; Fienberg, 1970; Franklin & Lorenz, 1989; Pukelsheim & Simeone, 2009; Qu et al., 2023). Notably, IPF solves a KL divergence minimization problem (Bregman, 1967; Ireland & Kullback, 1968; Csiszár, 1975; Ruschendorf, 1995), and is a coordinate descent type algorithm for its dual problem (Luo & Tseng, 1992). Recently, Qu et al. (2023) further established connections between IPF and ML estimation of choice models. We leverage these observations to identify a Poisson network model whose MLEs are recovered by IPF (Section 3). In Appendix B.2, we discuss other settings where IPF is known to recover MLEs of distinct models, such as contingency tables (Bishop & Fienberg, 1969; Bishop et al., 1974; Little & Wu, 1991; Little, 1993). The connections of IPF to Poisson type models also have precedence in economics, such as trade (Silva & Tenreyro, 2006) and matching (Galichon & Salanié, 2022a, b).
@@
Network inference. The problem of inferring dynamic networks from marginals has been central to the impactful guidance of COVID-19 policy from mobility data (Chang et al., 2021a, b; Chaudhuri et al., 2022; Li et al., 2023; Alimohammadi et al., 2023), and also appears across many domains, including transportation (Carey et al., 1981), communication (Kruithof, 1937), and migration (Plane, 1982; Pham & Komiyama, 2022). Prior works have also explored related but distinct network inference problems. When node-level signals are observed but the network is not known at all, temporal or spatial relationships are often used to infer which nodes are likely to be connected (Gomez-Rodriguez et al., 2012; Hallac et al., 2017; Akagi et al., 2018; Chen et al., 2021; Rossi et al., 2022). A closer setting to ours is where the structure of the graph and the node-level marginals are known (Kumar et al., 2015; Maystre & Grossglauser, 2017). Our setting is distinct since we study dynamic networks and have access to the time-aggregated network, not just the binary structure. Nevertheless, we note similarities, as both settings solve matrix balancing problems, with the ChoiceRank algorithm of Maystre & Grossglauser (2017) closely connected to IPF as well (Qu et al., 2023).
@@ 3 Biproportional Poisson Model: A Statistical Framework of IPF for Inferring Networks
Matrix balancing and IPF. IPF is an iterative algorithm that seeks to solve the following matrix balancing problem:
Given positive vectors $p\in\mathbb{R}_{++}^{m},q\in\mathbb{R}_{++}^{n}$ with $\sum p_{i}=\sum q_{j}$ and an initial non-negative matrix $X\in\mathbb{R}_{+}^{m\times n}$, find positive diagonal matrices $D^{0}$, $D^{1}$ satisfying the marginal conditions $D^{0}XD^{1}\cdot{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{1}_{n}}}}=p$ and $D^{1}X^{T}D^{0}\cdot{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{1}_{m}}}}=q$.
@@
subject to $\hat{Y}_{ij}\geq 0$, $\hat{Y}\mathbf{1}_{n}=p$, and $\hat{Y}^{T}\mathbf{1}_{m}=q$. (2) is feasible and bounded if and only if there exists $\hat{Y}$ with the desired marginals $p$ and $q$ and $\hat{Y}_{ij}=0$ whenever $X_{ij}=0$. Furthermore, the dual problem of (2) minimizes the potential function
@@
which is jointly convex in the dual variables $u\in\mathbb{R}^{m},v\in\mathbb{R}^{n}$. IPF is a coordinate descent type algorithm for (3) and $(d^{0},d^{1})$ is a solution to the matrix balancing problem if and only if $u=\log d^{0},v=-\log d^{1}$ is a minimizer of (3) (Luo & Tseng, 1992). For completeness, we provide details of the duality result in Appendix B.1.
Dynamic network inference problem. In our setting, we have a dynamic network with discrete time steps, where $X^{(t)}\in\mathbb{R}^{m\times n}_{+}$ represents the weighted adjacency matrix at time $t$. The network could be bipartite or unipartite (in which case $m=n$). We do not have access to $X^{(t)}$ due to privacy or sampling constraints, but we observe its row sums $p^{(t)}:=X^{(t)}\mathbf{1}_{n}$ and column sums $q^{(t)}:=(X^{(t)})^{T}\mathbf{1}_{m}$, as well as a time-aggregated network, $\bar{X}:=\sum_{t=1}^{T}X^{(t)}$, for some large $T$. The goal of the dynamic network inference problem is to provide a reasonable estimate of $X^{(t)}$, given $p^{(t)}$, $q^{(t)}$, and $\bar{X}$. The correspondence between this problem and the matrix balancing problem is natural: we can treat $\bar{X}$ as the initial matrix, $p^{(t)}$ as the target row marginals, and $q^{(t)}$ as the target column marginals. IPF’s solution to the matrix balancing problem with $\bar{X}$, $p^{(t)}$, and $q^{(t)}$ then serves as an estimate of the hourly network $X^{(t)}$. However, how should we interpret this estimate and why is IPF a justified approach here?
Our proposed network model. We provide a statistical justification of IPF as a solution to the network inference problem by identifying a generative network model under which IPF in fact recovers the MLEs of the network parameters. Our model, which we term the biproportional Poisson model, is defined as follows with parameters $u,v$:
@@
Model (4) posits independent Poisson samples $X^{(t)}_{ij}$ with expected value $\lambda_{ij}=e^{u_{i}}\bar{X}_{ij}e^{-v_{j}}$ wherever $\bar{X}_{ij}>0$ (we suppress time-indexing on $u$ and $v$ to simplify notation). Note that the model is unique up to normalization, e.g., $(u,v)^{T}\mathbf{1}_{m+n}=0$, since adding a constant $c$ to $(u,v)$ yields the same model. Since $X^{(t)}$ is not observed, it is a priori not obvious if the model parameters can be recovered by maximum likelihood estimation. Our first result reveals that although only $\bar{X}$, $p^{(t)}$, and $q^{(t)}$ are observed, MLEs of $u,v$ are well-defined, since $p^{(t)}$ and $q^{(t)}$ form the sufficient statistics of model (4). More importantly, our result connects IPF to (4), by showing that its MLEs are exactly the IPF solution to the matrix balancing problem, with (3) equivalent to the negative log-likelihood.
@@
We prove Theorem 3.1 in Appendix B.2. The core of our result lies in identifying a model whose log-likelihood is equivalent to $-g(u,v)$, since IPF minimizes $g(u,v)$ in (3). Our main intuition comes from Qu et al. (2023), who recently established connections between IPF and choice modeling, and observed that (3) reduces to the maximum likelihood objective of a general class of choice models.
@@
for all $i,j$ where $\bar{X}_{ij}>0$. We also verify that the stationarity assumption approximately holds on real-world data (Appendix E.3.3), thus justifying the decoupled approach. As two additional results, in Appendix B.4, we prove necessary and sufficient conditions for IPF to recover $X^{(t)}$ exactly, and in Appendix B.5 we prove that among a larger class of generalized linear models, the Poisson model is the unique one where the MLE is the IPF solution — so from the perspective of IPF, our generative model is canonical. We also note connections to (pseudo-)Poisson maximum likelihood regression in economic models of trade (Silva & Tenreyro, 2006) and matching (Galichon & Salanié, 2022a, b).
@@
Implications of our result. Our model allows us to interpret IPF through the lens of a generative network model. Since, in relevant applications, $X^{(t)}_{ij}$ is often the number of visits from node $i$ to $j$, (4) is consistent with queuing theory where rare events are modeled with a Poisson process due to the memory-less property. In addition, in the Poisson parameters $e^{u_{i}}\bar{X}_{ij}e^{-v_{j}}$, $u_{i}$ can be interpreted as the emission intensity of node $i$, and $-v_{j}$ as the absorption intensity of node $j$. For example, in mobility networks between residential neighborhoods and public places (Chang et al., 2021a), $u$ captures when each neighborhood is likelier to go out (e.g., younger populations more at night) and $-v$ captures each place’s visit propensity (e.g., schools visited more during the day while bars visited more at night). Moreover, since $d^{0}_{i}\bar{X}_{ij}d^{1}_{j}$ from IPF corresponds to $e^{u_{i}}\bar{X}_{ij}e^{-v_{j}}$, we can interpret the matrix inferred by IPF as estimating the expected values of the network-generating process.
@@
Defining an explicit model also yields several advantages. First, the equivalence between the IPF solution and this model’s MLE enables us to analyze IPF estimates using tools from statistical theory. In Section 4, we develop bounds on the MLE’s estimation error and establish that finite MLEs exist with high probability. Second, our model clarifies previously implicit assumptions when using IPF to infer dynamic networks, such as the stationarity assumption (6). Making such assumptions explicit allows practitioners to evaluate how reasonable the assumptions are given their domain and data; for example, we test several model assumptions on real-world bikeshare data (Appendix E.3.3). Third, defining an explicit model reveals natural ways to extend the model, such as non-Poisson distributions or interaction terms between rows and columns. These extensions allow us to test IPF under model misspecification (Appendix E.1.3) and create future opportunities for studying how changes in the model map back to changes in IPF. Finally, the model enables new empirical analyses of IPF (Section 6), such as quantifying uncertainty in the parameter estimates and evaluating IPF’s estimation of the network parameters, instead of only evaluating IPF’s error on the marginals, which is how IPF tends to be evaluated (Lovelace et al., 2015).

### v2

@@ 1 Introduction
One approach to solving this problem is to repurpose the classic iterative proportional fitting (IPF) procedure (Deming & Stephan, 1940), also widely known as Sinkhorn’s algorithm (Sinkhorn, 1974). Given a matrix $X$, target row marginals $p$, and target column marginals $q$, IPF tries to find a biproportional scaling of $X$ to match the target marginals. IPF poses an attractive solution for the network inference problem: it matches the marginal constraints of the problem, is computationally lightweight and space efficient, and is supported by decades of literature devoted to its analysis (Bregman, 1967; Fienberg, 1970; Csiszár, 1975; Pukelsheim & Simeone, 2009; Marino & Gerolin, 2020; Léger, 2021; Carlier, 2022; Qu et al., 2023). Furthermore, prior works have shown empirical success with using IPF to infer networks (Liang et al., 2006; McCord et al., 2010; Chang et al., 2021a), enabling important applications such as modeling epidemic spread on mobility networks (Li et al., 2023; Chaudhuri et al., 2022; Alimohammadi et al., 2023) and supporting policymakers (Chang et al., 2021b).
@@ 2 Related Work
Iterative proportional fitting.
IPF (Deming & Stephan, 1940), also known as Sinkhorn’s algorithm, biproportional fitting, raking, and the RAS algorithm, has a long history across disciplines. IPF offers many empirical advantages: it is straightforward to implement (see Algorithm 1 in Appendix A), transparent, reproducible, space efficient, and computationally lightweight (Liang et al., 2006; Lomax & Norman, 2015; Lovelace et al., 2015). The algorithmic properties of IPF have also been extensively studied (Sinkhorn, 1964; Fienberg, 1970; Franklin & Lorenz, 1989; Pukelsheim & Simeone, 2009; Qu et al., 2023). Notably, IPF solves a KL divergence minimization problem (Bregman, 1967; Ireland & Kullback, 1968; Csiszár, 1975; Ruschendorf, 1995), and is a coordinate descent type algorithm for its dual problem (Luo & Tseng, 1992). Recently, Qu et al. (2023) further established connections between IPF and ML estimation of choice models. We leverage these observations to identify a Poisson network model whose MLEs are recovered by IPF (Section 3). In Appendix B.3, we discuss other settings where IPF is known to recover MLEs of distinct models, such as contingency tables (Bishop & Fienberg, 1969; Bishop et al., 1974; Little & Wu, 1991; Little, 1993). The connections of IPF to Poisson-type models also have precedence in economics, such as trade (Silva & Tenreyro, 2006) and matching (Galichon & Salanié, 2022a, b). IPF (or Sinkhorn’s algorithm) is also closely connected to (discrete) entropy regularized optimal transport and Schrödinger bridges, which can be reformulated as matrix balancing problems with a kernel/reference matrix and fixed marginals, and solved using the IPF procedure (Cuturi, 2013; Marino & Gerolin, 2020; De Bortoli et al., 2021).
@@
Network inference.
The problem of inferring dynamic networks from marginals has been central to the impactful guidance of COVID-19 policy from mobility data (Chang et al., 2021a, b; Chaudhuri et al., 2022; Li et al., 2023; Alimohammadi et al., 2023), and also appears across many domains, including transportation (Carey et al., 1981), communication (Kruithof, 1937), and migration (Plane, 1982; Pham & Komiyama, 2022). Prior works have also explored related but distinct network inference problems. When node-level signals are observed but the network is not known at all, temporal or spatial relationships are often used to infer which nodes are likely to be connected (Gomez-Rodriguez et al., 2012; Hallac et al., 2017; Akagi et al., 2018; Chen et al., 2021; Rossi et al., 2022). A closer setting to ours is where the structure of the graph and the node-level marginals are known (Kumar et al., 2015; Maystre & Grossglauser, 2017). Our setting is distinct since we study dynamic networks and have access to the time-aggregated network, not just the binary structure. Nevertheless, we note similarities, as both settings solve matrix balancing problems, with the ChoiceRank algorithm of Maystre & Grossglauser (2017) closely connected to IPF as well (Qu et al., 2023).
Our network inference problem is also related to collective graphical models (CGMs), which aim to fit a model of individual behavior from aggregate data (Sheldon & Dietterich, 2011). Like CGMs, we also seek to estimate finer-grained information based on coarser data, but our setting differs in a number of ways: CGMs are typically applied in settings where only the time-varying marginals are known, such as to estimate population flows from counts per region over time (Iwata et al., 2017; Akagi et al., 2018; Iwata & Shimizu, 2019); CGMs try to learn a model of individual behavior, which is not our goal; and their mathematical formulations and estimation procedure are distinct from IPF. We compare our work to CGMs in greater detail in Appendix B.3.
@@ 3 Biproportional Poisson Model: A Statistical Framework of IPF for Inferring Networks
Matrix balancing and IPF.
IPF is an iterative algorithm that seeks to solve the following matrix balancing problem:
Given positive vectors $p\in\mathbb{R}_{++}^{m},q\in\mathbb{R}_{++}^{n}$ with $\sum_{i}p_{i}=\sum_{j}q_{j}=c$ and an initial non-negative matrix $X\in\mathbb{R}_{+}^{m\times n}$, find positive diagonal matrices $D^{0}$, $D^{1}$ satisfying the marginal conditions $D^{0}XD^{1}\cdot{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{1}_{n}}}}=p$ and $D^{1}X^{T}D^{0}\cdot{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{1}_{m}}}}=q$.
@@
subject to $\hat{Y}_{ij}\geq 0$, $\hat{Y}\mathbf{1}_{n}=p$, and $\hat{Y}^{T}\mathbf{1}_{m}=q$ [^2] Technically, the KL divergence $D_{\text{KL}}(P\|Q)$ is well-defined only on probability distributions $P,Q$. Without loss of generality, we may rescale $X$ so that $\sum_{ij}X_{ij}=\sum_{i}p_{i}=\sum_{j}q_{j}=c$. Then normalizing $\hat{Y},X$ by the common constant $c$ yields valid discrete probability distributions. The resulting KL minimization problem is equivalent to (2). (2) is feasible and bounded if and only if there exists $\hat{Y}$ with the desired marginals $p$ and $q$ and $\hat{Y}_{ij}=0$ whenever $X_{ij}=0$. Furthermore, the dual problem of (2) minimizes the potential function
@@
which is jointly convex in the dual variables $u\in\mathbb{R}^{m},v\in\mathbb{R}^{n}$. IPF is a coordinate descent type algorithm for (3) and $(d^{0},d^{1})$ is a solution to the matrix balancing problem if and only if $u=\log d^{0},v=-\log d^{1}$ is a minimizer of (3) (Luo & Tseng, 1992). For completeness, we provide details of the duality result in Appendix B.1. The KL minimization problem (2) is also closely related to entropy regularized optimal transport, which we discuss in more detail in Section B.2.
Dynamic network inference problem.
In our setting, we have a dynamic network with discrete time steps, where $X^{(t)}\in\mathbb{R}^{m\times n}_{+}$ represents the weighted adjacency matrix at time $t$. We do not have access to $X^{(t)}$ due to privacy or sampling constraints, but we observe its row sums $p^{(t)}:=X^{(t)}\mathbf{1}_{n}$ and column sums $q^{(t)}:=(X^{(t)})^{T}\mathbf{1}_{m}$, as well as a time-aggregated network, $\bar{X}:=\sum_{t=1}^{T}X^{(t)}$, for some large $T$. The goal of the dynamic network inference problem is to provide a reasonable estimate of $X^{(t)}$, given $p^{(t)}$, $q^{(t)}$, and $\bar{X}$. The correspondence between this problem and the matrix balancing problem is natural: we can treat $\bar{X}$ as the initial matrix, $p^{(t)}$ as the target row marginals, and $q^{(t)}$ as the target column marginals. IPF’s solution to this matrix balancing problem then serves as an estimate of the hourly network $X^{(t)}$. However, how should we interpret this estimate and why is IPF a justified approach here?
Our proposed network model.
We provide a statistical justification of IPF as a solution to the network inference problem by identifying a generative network model under which IPF in fact recovers the MLEs of the network parameters. Our model, which we term the biproportional Poisson model, is defined as follows with parameters $u,v$:
@@
Model (4) posits independent Poisson samples $X^{(t)}_{ij}$ with expected value $\lambda_{ij}=e^{u_{i}}\bar{X}_{ij}e^{-v_{j}}$ wherever $\bar{X}_{ij}>0$ (we suppress time-indexing on $u$ and $v$ to simplify notation). Note that the model is unique up to normalization, e.g., $(u,v)^{T}\mathbf{1}_{m+n}=0$, since adding a constant $c$ to $(u,v)$ yields the same model. Since $X^{(t)}$ is not observed, it is a priori not obvious if the model parameters can be recovered by maximum likelihood estimation. Our first result reveals that although only $\bar{X}$, $p^{(t)}$, and $q^{(t)}$ are observed, MLEs of $u,v$ are well-defined, since $p^{(t)}$ and $q^{(t)}$ form the sufficient statistics of model (4). More importantly, our result connects IPF to (4), by showing that its MLEs are exactly the IPF solution to the matrix balancing problem, with (3) equivalent to the negative log-likelihood. This connection also partially inspired the name “biproportional Poisson”, as an alternative name of IPF is “biproportional fitting” (Bacharach, 1965).
@@
We prove Theorem 3.1 in Appendix B.3. The core of our result lies in identifying a model whose log-likelihood is equivalent to $-g(u,v)$, since IPF minimizes $g(u,v)$ in (3). Our result is closely related to Qu et al. (2023), who recently established connections between IPF and choice modeling, and observed that (3) reduces to the maximum likelihood objective of a general class of choice models.
@@
for all $i,j$ where $\bar{X}_{ij}>0$. We also verify that the stationarity assumption approximately holds on real-world data (Appendix E.3.3), thus justifying the decoupled approach. As two additional results, in Appendix B.5, we prove necessary and sufficient conditions for IPF to recover $X^{(t)}$ exactly, and in Appendix B.6, we prove that among a larger class of generalized linear models, the Poisson model is the unique one where the MLE is the IPF solution — so from the perspective of IPF, our generative model is canonical.
Implications of our result.
@@
Our model allows us to interpret IPF through the lens of a generative network model. Since, in relevant applications, $X^{(t)}_{ij}$ is often the number of visits from node $i$ to $j$, (4) is consistent with queuing theory where rare events are modeled with a Poisson process due to the memory-less property. In addition, in the Poisson parameters $e^{u_{i}}\bar{X}_{ij}e^{-v_{j}}$, $u_{i}$ can be interpreted as the emission intensity of node $i$, and $-v_{j}$ as the absorption intensity of node $j$. For example, in mobility networks between residential neighborhoods and public places (Chang et al., 2021a), $u$ captures when each neighborhood is likelier to go out (e.g., younger populations more at night) and $-v$ captures each place’s visit propensity (e.g., schools visited more during the day while bars visited more at night). Notably, the biproportional form of the model assumes that there are not time-varying interactions between rows and columns (e.g., if a place offers special discounts for seniors at this time, attracting neighborhoods with large senior populations), making explicit one of the key assumptions of using IPF in this network inference setting. Moreover, since $d^{0}_{i}\bar{X}_{ij}d^{1}_{j}$ from IPF corresponds to $e^{u_{i}}\bar{X}_{ij}e^{-v_{j}}$, we can interpret the matrix inferred by IPF as estimating the expected values of the network-generating process.
We also note connections of our model to (pseudo-)Poisson maximum likelihood regression in economics (Gourieroux et al., 1984), including for models of trade (Silva & Tenreyro, 2006) and matching (Galichon & Salanié, 2022a, b). In particular, Theorem 3.1 implies that estimating the biproportional model is equivalent to solving a pseudo-Poisson maximum likelihood problem. Consequently, the results of Gourieroux et al. (1984) guarantee that the minimizers of Equation 3 are consistent estimators of $u,v$ even under misspecifications of the distribution and homoskedasticity of $X^{(t)}_{ij}$ given $\lambda_{ij}=e^{u_{i}}\bar{X}_{ij}e^{-v_{j}}$, highlighting the robustness of our biproportional Poisson model.
@@
Defining an explicit model also yields several advantages. First, the equivalence between the IPF solution and this model’s MLE enables us to analyze IPF estimates using tools from statistical theory. In Section 4, we develop bounds on the MLE’s estimation error and establish that finite MLEs exist with high probability. Second, our model clarifies previously implicit assumptions when using IPF to infer dynamic networks, such as the stationarity assumption (6) or the lack of time-varying interactions. Making such assumptions explicit allows practitioners to evaluate how reasonable the assumptions are given their domain and data; for example, we test several model assumptions on real-world bikeshare data (Appendix E.3.3). Third, defining an explicit model reveals natural ways to extend the model, such as non-Poisson distributions or interaction terms between rows and columns. These extensions allow us to test IPF under model misspecification (Appendix E.1.3) and create future opportunities for studying how changes in the model map back to changes in IPF. Finally, the model enables new empirical analyses of IPF (Section 6), such as quantifying uncertainty in the parameter estimates and evaluating IPF’s estimation of the network parameters, instead of only evaluating IPF’s error on the marginals, which is how IPF tends to be evaluated (Lovelace et al., 2015).